(Efficiency analysis)  The Brenmar Sales Company had a gross profit margin​ (gross

profits divided by÷​sales)

26 percent and sales of

$ 9.9 million last year.  

78 percent of the​ firm's sales are on​ credit, and the remainder are cash sales. ​ Brenmar's current assets equal

$1.8 million, its current liabilities equal

$ 300,700, and it has

$100,100

in cash plus marketable securities.a. If​ Brenmar's accounts receivable equal $563,200​,

what is its average collection​ period?b. If Brenmar reduces its average collection period to 20

​days, what will be its new level of accounts​ receivable?c.  ​Brenmar's inventory turnover ratio is 8.2

times. What is the level of​ Brenmar's inventories?

a.  If​ Brenmar's accounts receivable equal $563,200​,

what is its average collection​ period?The​ company's average collection period will be

nothing

days.  ​(Round to two decimal​ places.)
Answer #1
Sales 9,900,000
Cogs 9,900,000 * 74% = 7,326,000
Gross Profit 9,900,000 * 26% = 2,574,000
Credit sales 9,900,000 * 78% = 7,722,000
A
Account receivable 563,200
Receivable turnover ratio Credit sales / Accounts receivables 7,722,000 / 563,200 = 13.71
Average collection period 365 / Receivable turnover ratio 365/13.71 = 27 days
B
Average collection period reduces to 20 days
Receivable turnover ratio 365 / average collection period 365/20 = 18.25
Account receivable Credit sales / Receivable turnover ratio 7,722,000/ 18.25 =423,123.29
C
Inventory Cogs / Inventory turnover ratio 7,326,000/8.2 = 893,414.63